Im so Lost could u please explain to me how to fin the sum 1+2....+80
Answer by fcabanski(1391)

This is a series (sum of) an arithmetic sequence. The common difference is 1. Each term is 1 more than the last term.


This sequence could also have other rules, such as "the first two terms are 1 and 2, and then all other terms are 80." But for this series problem, we have to assume it's an arithmetic sequence.


The sum of a sequence (a series) 1+2+...n is S(n) = (n(a(1)+a(n))/2 where a(1) is the first term, a(n) is the nth term, n is the number in question (i.e. the desired term #), and S(n) is the sum of the terms.


S(n) = 80(1+80)/2 = 80*81/2 = 40*81 = 3240